# Kinetic theory of point vortices at order ${1/N}$ and ${1/N^{2}}$

### v1

@@ Appendix G Numerical simulations
Bootstraps. In practice, it is essential to estimate the uncertainties associated with these measurements. To that purpose, we use bootstraps. A given bootstrap proceeds as follows. For every value of ${(N,q)}$, we perform the ensemble average of ${N(\!<\!J,t)}$ over a sample of $N_{\mathrm{real}}$ realisations drawn, with repetitions, from the $N_{\mathrm{real}}$ realisations available. Using the associated ${\langle N(\!<\!J,t)\rangle}$, we perform the linear fit from Eq. (69) from which we also estimate the variance of the slope $\beta$. Finally, for each ${(N,q)}$ and each considered action, we can draw a value of $\beta$ according to the associated Gaussian distribution. In practice, we performed a total of 1 000 bootstrap measurements, with which we could obtain the 16% and 84% level contours presented in Figs. 5 and 12, as well as the median values presented in Figs. 7 and 13.
Appendix H Heuristic estimate of the diffusion coefficient
When derived from the BBGKY hierarchy at order ${1/N}$, the Landau equation generically reads Chavanis (2013)
$\frac{\partial F}{\partial t}={}$ $\frac{\partial}{\partial J}\bigg{[}\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}\tau\,\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\,\!\!\int_{0}^{2\pi}\!\!\!\mathrm{d}\theta_{1}\,$
×$\times\big{\langle}\mathcal{F}(1\rightarrow 0,t)\,G(t,t-\tau)\,\mathcal{F}(1\rightarrow 0,t-\tau)\big{\rangle}$
×$\times\bigg{(}\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg{)}F(J,t)\,\frac{1}{\gamma}F(J_{1},t)\bigg{]},$ (70)
with
$\mathcal{F}(1\rightarrow 0,s)=-\mathrm{i}\gamma\sum_{k}k\,\mathrm{e}^{\mathrm{i}k[\theta(s)-\theta_{1}(s)]}\,U_{k}(J,J_{1}),$ (71)
the force (velocity) created by particle (vortex) $1$ on particle (vortex) 00 at time $s$, and ${G(t,t\!-\!\tau)}$ is an appropriate Green function taking into account the motion of the particles between $t$ and ${t\!-\!\tau}$ Chavanis (2013). In Eq. (70), the brackets indicate an average over the fluctuating trajectories of the particles (see below). To calculate the temporal correlation function in Eq. (70), one usually assumes that the particles follow the unperturbed mean field trajectories at leading order. This leads to the Landau equation (Eq. 14). However, this approximation breaks down close to an extremum of ${\Omega(J)}$, i.e. when particles have similar frequencies. In that case, one has to take into account fluctuations in the particles’ individual trajectories.
Following Dubin (2003), let us consider a simplified model where we approximate the motion of the particles with the effective stochastic equations
$\frac{\mathrm{d}J}{\mathrm{d}t}=0,\qquad\frac{\mathrm{d}\theta}{\mathrm{d}t}=\Omega(J)+\sqrt{2D_{\theta\theta}}\,\eta(t),$ (72)
where ${\eta(t)}$ is a Gaussian white noise such that ${\langle\eta(t)\rangle\!=\!0}$ and ${\langle\eta(t)\,\eta(t^{\prime})\rangle\!=\!\delta_{\mathrm{D}}(t\!-\!t^{\prime})}$, and $D_{\theta\theta}$ is a diffusion coefficient in angles. Integrating these equations between $t$ and ${t\!-\!\tau}$, substituting the stochastic trajectories ${J(t\!-\!\tau)\!=\!J}$ and ${\theta(t\!-\!\tau)\!=\!\theta\!-\!\Omega(J)\tau\!-\!\!\int_{0}^{\tau}\!\mathrm{d}t^{\prime}\eta(t\!-\!t^{\prime})}$ into Eq. (70), and carrying out the integration over $\theta_{1}$, we obtain
$\frac{\partial F}{\partial t}={}$ $4\pi\gamma\frac{\partial}{\partial J}\bigg{[}\!\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}J_{1}\,\!\!\int_{0}^{+\infty}\!\!\!\!\mathrm{d}\tau\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$
@@ Appendix H Heuristic estimate of the diffusion coefficient
×$\times\mathrm{e}^{\mathrm{i}k\Delta\Omega(J,J_{1})\tau}\left\langle\mathrm{e}^{\mathrm{i}k\!\int_{0}^{\tau}\!\mathrm{d}t^{\prime}[\eta(t-t^{\prime})-\eta_{1}(t-t^{\prime})]}\right\rangle$
×$\times\bigg{(}\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg{)}F(J,t)F(J_{1},t)\bigg{]},$ (73)
where we introduced the frequency difference, ${\Delta\Omega(J,J_{1})\!=\!\Omega(J)\!-\!\Omega(J_{1})}$. If we assume a Gaussian distribution for the fluctuations Dubin (2003), we have

### v2

@@ Appendix G Numerical simulations
Bootstraps. In practice, it is essential to estimate the uncertainties associated with these measurements. To that purpose, we use bootstraps. A given bootstrap proceeds as follows. For every value of (N,q){(N,q)}, we perform the ensemble average of N(<J,t){N(\!<\!J,t)} over a sample of $N_{\mathrm{real}}$ realisations drawn, with repetitions, from the $N_{\mathrm{real}}$ realisations available. Using the associated $\langle N(\!<\!J,t)\rangle$}, we perform the linear fit from Eq. (78) from which we also estimate the variance of the slope $\beta$. Finally, for each (N,q){(N,q)} and each considered action, we can draw a value of $\beta$ according to the associated Gaussian distribution. In practice, we performed a total of 1 000 bootstrap measurements, with which we could obtain the 16% and 84% level contours presented in Figs. 5 and 13, as well as the median values presented in Figs. 8 and 14.
Diffusion coefficients. In order to measure the diffusion coefficients in Figs. 5, 7, we proceeded as follows. We start from the exact same setup as in the previous $N$-body runs, using N=2 000{N\!=\!2\,000} and q=0.0001{q\!=\!0.0001}. We add to each simulation 1 000 zero-mass test particles, which, therefore, do not contribute the system’s instantaneous potential. Initially, these particles are all sampled with the same initial action, $J_{\mathrm{test}}$}, and with their initial angle, $\theta$, drawn uniformly within $0,2\pi]$}. In practice, we considered sets of initial conditions with $0.8\!\leq\!J_{\mathrm{test}}/J_{0}\!\leq\!1.2$} in increments of $10^{-3}\,J_{0}$. For each such choice of $J_{\mathrm{test}}$, we performed 512 independent realisations. Finally, using the same integration parameters as previously (timestep, number of dumps, etc.), for each realisation, we record the time series of $t\!\mapsto\!\langle(J(t)\!-\!J(0))^{2}\rangle$}, averaged over all the test particles from that realisation. In Fig. 18, we illustrate such a time series, once averaged over all available realisations.
Figure 18: Time evolution of the (squared) change in action, $\langle\Delta J^{2}(t)\rangle$}, of test particles in numerical simulations, when averaged over all available realisations, for profile (1). Different colours correspond to different initial actions, $J_{\mathrm{test}}$, for the test particles. The diffusion of the test particles shows the transition from a ballistic regime ($\propto\!t^{2}$}) at early times to a diffusive one ($\propto\!t$}) at late times.
In that figure, we note a transition between an initial “ballistic” regime, during which $\langle\Delta J^{2}(t)\rangle$} grows like $t^{2}$, and a “diffusive” regime during which it only grows like t{t}. In Fig. 18, we find that the typical relative change in $\langle\Delta J^{2}\rangle/J_{0}^{2}$} at the end of the integration is, at minimum, of order $10^{-7}$. This is much larger than $10^{-13}$, the typical relative error in the $N$-body invariants (see Fig. 16). This brings some confidence in the appropriate convergence of the $N$-body measurements of the diffusion coefficients.
In order to measure the diffusion coefficient, we estimate the numerical diffusion coefficient, $D_{\mathrm{num}}$, through the slope
$D_{\mathrm{num}}(J_{\mathrm{test}})=\frac{\langle\Delta J^{2}(t_{\mathrm{max}})\rangle-\langle\Delta J^{2}(t_{\mathrm{min}})\rangle}{t_{\mathrm{max}}-t_{\mathrm{max}}},$ (79)
between the times $t_{\mathrm{min}}$ and $t_{\mathrm{max}}$. For Fig. 5, in order not to be polluted by the initial ballistic regime (Fig. 18), we considered the time range $t_{\mathrm{min}}\!=\!500\,T_{\mathrm{dyn}}$} and $t_{\mathrm{max}}\!=\!1\,000\,T_{\mathrm{dyn}}$}. We checked that varying slightly these values did not lead to any significant change in the measured $D_{\mathrm{num}}$. For the finite-time diffusion coefficients presented in Fig. 7, we fixed $t_{\mathrm{min}}\!=\!0$} and used the values of $t_{\mathrm{max}}$ indicated in that figure. In both figures, uncertainties are estimated using bootstrap resamplings. This amounts to performing 1 000 different ensembles averages over sets of 512 realisations, but allowing for repetitions. In Figs. 5 and 7, we present the 16% and 84% level contours over these bootstraps.
Appendix H Properties of the regularised Landau equation
In this Appendix, we study the properties of the regularised Landau equation from Eq. (32). In particular, we show how resonance broadening regularises the divergences of the Landau diffusion coefficient for $2D$ point vortices.
H.1 Regularisation of the logarithmic divergence of the Landau diffusion coefficient
Let us first consider the case of a monotonic frequency profile, or a value of $J$ far from the value $J_{\star}$, where the frequency profile has an extremum.
The regularised diffusion coefficient is given by Eq. (33). Using the expression of the unsoftened potential of interaction from Eq. (52) and the identity $\sum_{k>0}x^{k}/k\!=\!-\ln(1\!-\!x)$}, we can explicitly perform the sum over $k$ in the function $\Lambda_{\mathrm{tot}}(J,J_{1})$} Chavanis (2012a), and obtain
$D(J)=-\tfrac{1}{2}\gamma\!\!\int_{0}^{+\infty}\!\!\!\!\!\!\mathrm{d}J_{1}\,\ln\left(1-\frac{J_{\mathrm{min}}}{J_{\mathrm{max}}}\right)$
$\times\,\delta_{T_{\mathrm{reg}}}\left[\Omega(J)-\Omega(J_{1})\right]F(J_{1}),$ (80)
with $J_{\mathrm{min}}\!=\!\mathrm{Min}[J,J_{1}]$} and similarly for $J_{\mathrm{max}}$.
When $T_{\mathrm{reg}}\!\rightarrow\!+\infty$}, the Lorentzian tends to a Dirac delta imposing $\Omega(J_{1})\!=\!\Omega(J)$}, hence $J_{1}\!=\!J$}. In that case, the (Landau) diffusion coefficient diverges logarithmically at small scales, i.e. for $k\!\rightarrow\!+\infty$}. Introducing a cut-off $k_{\mathrm{max}}$, and using Eq. (23), we obtain Chavanis (2012a)
$D(J)=\tfrac{1}{2}\gamma\frac{F(J)}{|\Omega^{\prime}(J)|}\ln\Lambda,$ (81)
with $\ln\Lambda\!=\!\sum_{k>0}\frac{1}{k}\sim\ln k_{\mathrm{max}}$}. When $T_{\mathrm{reg}}$ is finite, there is no divergence in Eq. (80).
Let us now assume that $T_{\mathrm{reg}}$ is sufficiently large. In that case, the regularised Dirac delta (Lorentzian) is strongly peaked around $J_{1}\!=\!J$}. We can therefore make a local approximation and expand $\Omega(J_{1})$} to first order around $J$. We can also replace $F(J_{1})$} by F(J){F(J)}. Doing so, at leading order, we obtain
@@ H.1 Regularisation of the logarithmic divergence of the Landau diffusion coefficient
$D(J)\sim-\tfrac{1}{2}\gamma F(J)J\!\!\int_{-\infty}^{+\infty}\!\!\!\!\!\!\mathrm{d}x\,\ln|x|\,\delta_{T_{\mathrm{reg}}}\left[\Omega^{\prime}(J)Jx\right],$ (82)
giving
$D(J)\sim\tfrac{1}{2}\gamma\frac{F(J)}{|\Omega^{\prime}(J)|}\ln\left[T_{\mathrm{reg}}|\Omega^{\prime}(J)|J\right].$ (83)
We thus find that the regularized Coulomb logarithm, $\ln\Lambda$}, is equal to the logarithm of the ratio between $T_{\mathrm{reg}}$ and the timescale $1/(|\Omega^{\prime}(J)|J)$} set by the shear of the frequency profile.
H.2 Regularisation of the Landau diffusion coefficient when $J\!=\!J_{\star}$}
We now consider the case where the frequency profile has a minimum at $J_{\star}$, i.e. one has $\Omega^{\prime}_{\star}\!=\!\Omega^{\prime}(J_{\star})\!=\!0$}. In that case, the Landau diffusion coefficient diverges at $J\!=\!J_{\star}$}, as discussed in the main text (see also Eq. 81). Resonance broadening, which is here accounted for heuristically by a finite value of $T_{\mathrm{reg}}$, regularises this divergence. Let us consider the function $D_{\star}(T_{\mathrm{reg}})$} obtained by taking $J\!=\!J_{\star}$} in Eq. (80).
For $T_{\mathrm{reg}}\!\gg\!T_{\mathrm{reg}}^{\mathrm{c}}$} (with the transition timescale $T_{\mathrm{reg}}^{\mathrm{c}}$ to be introduced in Eq. 89), and using the same method as above but pushing the Taylor expansion up to second order (assuming $\Omega^{\prime\prime}_{\star}\!=\!\Omega^{\prime\prime}(J_{\star})\!\neq\!0$}), we find at leading order
$D_{\star}\sim-\tfrac{1}{2}\gamma F_{\star}J_{\star}\!\!\int_{-\infty}^{+\infty}\!\!\!\!\!\!\mathrm{d}x\,\ln|x|\,\delta_{T_{\mathrm{reg}}}\left[\tfrac{1}{2}\Omega_{\star}^{\prime\prime}J_{\star}^{2}x^{2}\right],$ (84)
giving
$D_{\star}\sim\tfrac{1}{8}\gamma F_{\star}\sqrt{\frac{T_{\mathrm{reg}}}{|\Omega_{\star}^{\prime\prime}|}}\left\{\pi+2\ln\left[\tfrac{1}{2}|\Omega_{\star}^{\prime\prime}|J_{\star}^{2}T_{\mathrm{reg}}\right]\right\}.$ (85)
Ignoring logarithmic corrections, and using $\Gamma\sim F_{\star}J_{\star}$, we obtain the scaling
$D_{\star}\sim\gamma\Gamma\sqrt{\frac{T_{\mathrm{reg}}}{J_{\star}^{2}\,|\Omega_{\star}^{\prime\prime}|}}\propto T_{\mathrm{reg}}^{1/2}.$ (86)
This scaling remains valid for a softened potential.
For $T_{\mathrm{reg}}\!\ll\!T_{\mathrm{reg}}^{\mathrm{c}}$}, using $\delta_{T_{\mathrm{reg}}}(\omega)\!\rightarrow\!T_{\mathrm{reg}}/\pi$}, we obtain
$D_{\star}\sim-\tfrac{1}{2}\gamma\frac{T_{\mathrm{reg}}}{\pi}\!\!\int_{0}^{+\infty}\!\!\!\!\!\!\mathrm{d}J_{1}\,\ln\left(1-\frac{J_{\mathrm{min}}}{J_{\mathrm{max}}}\right)F(J_{1}),$ (87)
yielding the scaling
$D_{\star}\sim\gamma\Gamma T_{\mathrm{reg}}\propto T_{\mathrm{reg}}.$ (88)
Once again, this scaling remains valid for a softened potential. The transition between these two regimes occurs at
$T_{\mathrm{reg}}^{\mathrm{c}}\sim\frac{1}{J_{\star}^{2}\,|\Omega_{\star}^{\prime\prime}|},\qquad D_{\star}^{\mathrm{c}}\sim\frac{\gamma\Gamma}{J_{\star}^{2}\,|\Omega_{\star}^{\prime\prime}|}.$ (89)
We note that the value of $T_{\mathrm{reg}}^{\mathrm{c}}$ is independent of $N$. Since $T_{\mathrm{reg}}$ increases with $N$, we expect that we are generically in the regime $T_{\mathrm{reg}}\!\gg\!T_{\mathrm{reg}}^{\mathrm{c}}$}, not in the regime $T_{\mathrm{reg}}\!\ll\!T_{\mathrm{reg}}^{\mathrm{c}}$}.
H.3 Self-consistent relation
In the main text, $T_{\mathrm{reg}}$ was treated as an adjustable parameter, see e.g., Fig. 6. Now, if we assume the relation
$T_{\mathrm{reg}}\sim\frac{J_{\star}^{2}}{D_{\star}},$ (90)
and use Eq. (80) with $J\!=\!J_{\star}$}, we can obtain the expression of $T_{\mathrm{reg}}$ (or $D_{\star}$) self-consistently. We refer to Appendix I for a graphical study.
For $T_{\mathrm{reg}}\!\gg\!T_{\mathrm{reg}}^{\mathrm{c}}$} (i.e. for $D_{\star}\!\gg\!D_{\star}^{\mathrm{c}}$}), combining Eqs. (86) and (90), we get
$D_{\star}\sim\left(\frac{\gamma\Gamma}{\sqrt{|\Omega_{\star}^{\prime\prime}|}}\right)^{2/3},\quad T_{\mathrm{reg}}\sim J_{\star}^{2}\left(\frac{\sqrt{|\Omega_{\star}^{\prime\prime}|}}{\gamma\Gamma}\right)^{2/3}.$ (91)
This gives $T_{\mathrm{reg}}\!\propto\!D_{\star}^{-1}\!\propto\!\gamma^{-2/3}\!\propto\!N^{2/3}$}. To the best of our knowledge, such a $N^{2/3}$ scaling has not been reported before.
For $T_{\mathrm{reg}}\!\ll\!T_{\mathrm{reg}}^{\mathrm{c}}$} (i.e. for $D_{\star}\!\ll\!D_{\star}^{\mathrm{c}}$}), combining Eqs. (88) and (90), we get
$D_{\star}\sim J_{\star}\sqrt{\gamma\Gamma},\quad T_{\mathrm{reg}}\sim\frac{J_{\star}}{\sqrt{\gamma\Gamma}},$ (92)
giving $T_{\mathrm{reg}}\!\sim\!D_{\star}^{-1}\!\sim\!\gamma^{-1/2}\!\sim\!N^{1/2}$}. This is the $N^{1/2}$ scaling presented in the main text, which implicitly assumed $T_{\mathrm{reg}}\!\ll\!T_{\mathrm{reg}}^{\mathrm{c}}$} (see Section III.5). It corresponds to the Taylor-McNamarra Taylor and McNamara (1971) scaling which applies to an unsheared flow (see Appendix I.3). In the case considered in Taylor and McNamara (1971), $\Omega(J)$} is constant, so that $\Omega^{\prime\prime}(J)$} vanishes everywhere, as well as all its higher order derivatives. This implies $T_{\mathrm{reg}}^{\mathrm{c}}\!\rightarrow\!+\infty$} and $D_{\star}^{\mathrm{c}}\!\rightarrow\!0$}, following Eq. (89). By contrast, when $\Omega_{\star}^{\prime\prime}\!\neq\!0$}, as in our case, there exists another regime, $T_{\mathrm{reg}}\!\gg\!T_{\mathrm{reg}}^{\mathrm{c}}$}, that yields a $N^{2/3}$ scaling. It could be that this scaling, which is valid when $|\Omega_{\star}^{\prime\prime}|\!\gg\!\sqrt{\gamma\Gamma}/J_{\star}^{3}\!\sim\!1/\sqrt{N}$}, is more relevant than the $N^{1/2}$ scaling, valid in the opposite limit, presented in the main text. Unfortunately, these two scalings cannot be distinguished in the numerical simulations presented here, since $2/3$ is close to $1/2$. This subtle distinction will be considered in a future work, by studying a simpler system for which more precise numerical simulations can be performed.
Remark. One can generalise the previous results to the case where the first n1{n\!-\!1} derivatives of $\Omega(J)$} vanish at $J_{\star}$ but $\Omega_{\star}^{(n)}\!=\!\Omega^{(n)}(J_{\star})\!\neq\!0$}. Instead of Eq. (86), we find
$D_{\star}\sim\frac{\gamma\Gamma T_{\mathrm{reg}}^{1-1/n}}{J_{\star}\big|\Omega_{\star}^{(n)}\big|^{1/n}}$ (93)
for $T_{\mathrm{reg}}\!\gg\!1/(J_{\star}^{n}|\Omega_{\star}^{(n)}|)$}. Then, instead of Eq. (91), we find
$T_{\mathrm{reg}}\sim\frac{J_{\star}^{2}}{D_{\star}}\sim\left[\frac{J_{\star}^{3}\big|\Omega_{\star}^{(n)}\big|^{1/n}}{\gamma\Gamma}\right]^{\frac{1}{2-1/n}}$ (94)
for $|\Omega_{\star}^{(n)}|\!\gg\!\sqrt{\gamma\Gamma}/J_{\star}^{n+1}\!\sim\!1/\sqrt{N}$}. This yields the general scaling $T_{\mathrm{reg}}\!\sim\!D_{\star}^{-1}\!\sim\!\gamma^{-1/(2-1/n)}\!\sim\!N^{1/(2-1/n)}$}. The results from Eqs. (87) and (92), corresponding to the Taylor–McNamara regime (Taylor and McNamara, 1971), are not changed. For n=1{n\!=\!1} we recover the Landau scaling, for n=2{n\!=\!2} we recover the new scaling from Eq. (91), and for $n\!\rightarrow\!+\infty$} we recover the Taylor–McNamara (Taylor and McNamara, 1971) scaling. Exploring these different regimes of relaxation will be the topic of a future work.
Appendix I Resonance broadening from stochastic equations of motion
In this Appendix, we introduce an alternative stochastic model to justify the heuristic resonance broadening from Section III.5 and motivate the regularisation of the Dirac delta by a Lorentzian function (Eq. 31). In addition, we use this model to estimate the diffusion coefficient of point vortices. We also point out the limitations of this approach and the need for a more refined kinetic framework. Finally, we discuss the connection between this work and previous literature.
I.1 Regularised Landau equation from individual stochastic equations
When derived from the BBGKY hierarchy at order 1/N{1/N}, the Landau equation generically reads Chavanis (2013)
$\frac{\partial F(J,t)}{\partial t}={}$ $\frac{\partial}{\partial J}\bigg[\!\int_{0}^{+\infty}\!\!\!\!\!\!\mathrm{d}\tau\,\!\int_{0}^{+\infty}\!\!\!\!\!\!\mathrm{d}J_{1}\,\!\!\int_{0}^{2\pi}\!\!\!\!\mathrm{d}\theta_{1}\,$
$\times\big\langle\mathcal{F}(1\rightarrow 0,t)\,G(t,t-\tau)\,\mathcal{F}(1\rightarrow 0,t-\tau)\big\rangle$
$\times\bigg(\frac{\partial}{\partial J}-\frac{\partial}{\partial J_{1}}\bigg)F(J,t)\,\frac{1}{\gamma}F(J_{1},t)\bigg],$ (95)
where
$\mathcal{F}(1\rightarrow 0,s)=-\mathrm{i}\gamma\sum_{k}k\,\mathrm{e}^{\mathrm{i}k[\theta(s)-\theta_{1}(s)]}\,U_{k}(J,J_{1})$ (96)
is the force (velocity) created by particle (vortex) $1$ on particle (vortex) 0 at time $s$, and $G(t,t\!-\!\tau)$} is an appropriate Green function taking into account the motion of the particles between $t$ and $t\!-\!\tau$} Chavanis (2013). In Eq. (95), the brackets indicate an average over the fluctuating trajectories of the particles (see below). To calculate the temporal correlation function in Eq. (95), one usually assumes that the particles follow the unperturbed mean field trajectories at leading order. This leads to the Landau equation (Eq. 14). However, this approximation breaks down notably close to an extremum of $\Omega(J)$}, i.e. when particles have similar frequencies. In that case, one has to take into account fluctuations in the particles’ individual trajectories.
Following Dubin (2003), we replace the deterministic equations of motion by effective stochastic equations where the amplitude of the noise term is determined by the diffusion coefficient obtained from the kinetic theory itself (see below). Assuming that the diffusion is isotropic Dubin (2003), we then have $D_{\theta\theta}(J)\!=\!D(J)/(4J^{2})$}, where $D(J)\!\equiv\!D_{JJ}(J)$} [resp. $D_{\theta\theta}(J)$}] is the diffusion coefficient in action [resp. angle].
Since our goal is to illustrate the origin of resonance broadening in a simple manner, we make the following approximations: (i) We assume that the stochastic process can be described by a Gaussian white noise; (ii) We ignore the diffusion in action in the stochastic equations of motion; (iii) We assume that the diffusion coefficient in angle, $D_{\theta\theta}$, is constant with value $D_{\star}/(4J_{\star}^{2})$}, where $D_{\star}\!=\!D(J_{\star})$} is the action diffusion coefficient at the extremum of the frequency profile. As a result, we approximate the motion of the particles with the effective stochastic equations
$\frac{\mathrm{d}J}{\mathrm{d}t}{}$ $=0,$ (97a)
$\frac{\mathrm{d}\theta}{\mathrm{d}t}{}$ $=\Omega(J)+\sqrt{D_{\star}/(4J_{\star}^{2})}\,\eta(t),$ (97b)
where $\eta(t)$} is a Gaussian white noise such that $\langle\eta(t)\rangle\!=\!0$} and $\langle\eta(t)\,\eta(t^{\prime})\rangle\!=\!\delta_{\mathrm{D}}(t\!-\!t^{\prime})$}. Integrating these equations between $t$ and $t\!-\!\tau$}, we obtain the stochastic trajectories
$J(t\!-\!\tau){}$ $=J,$ (98a)
$\theta(t\!-\!\tau){}$ $=\theta\!-\!\Omega(J)\tau\!-\!\sqrt{D_{\star}/(4J_{\star}^{2})}\!\int_{0}^{\tau}\!\!\mathrm{d}t^{\prime}\,\eta(t\!-\!t^{\prime}).$ (98b)
Substituting these solutions into Eq. (95) and carrying out the integration over $\theta_{1}$, we obtain
$\frac{\partial F(J,t)}{\partial t}=4\pi\gamma\frac{\partial}{\partial J}\bigg[\!\!\int_{0}^{+\infty}\!\!\!\!\!\!\!\!\!\mathrm{d}J_{1}\,\!\!\int_{0}^{+\infty}\!\!\!\!\!\!\!\!\!\mathrm{d}\tau\sum_{k>0}k^{2}\,|U_{k}(J,J_{1})|^{2}$